OpenStudy (anonymous):

Find the slope of the line through the given points. (sqrt(2),-8), (2,-9) Would the answer be 1/sqrt(2)-2 ?

OpenStudy (anonymous):

the points (x1,y1) and (x2,y2) slope=(y2-y1)/(x2-x1) here slope=[-9-(-8)]/[2-sqrt(2)]=-1/[2-sqrt(2)]=1/[sqrt(2)-2] your answer is correct.